# Vistoria: A Multimodal System to Support Fictional Story Writing through Instrumental Text–Image Co-Editing

### v2

@@ 4. Vistoria System
In this section, we present the key features of Vistoria. As shown in Figure 2, the interface comprises three primary components: a left text editor, a central collapsible cluster panel, and a right canvas interface. The text editor houses the current story draft, serving as contextual information for content generation. The right canvas supports freeform sketching, text input, and comprehensive editing tools. The central cluster panel aggregates highlights and annotations from canvas content, organizing them by characters, objects, and scenes for reference and summary.
Together, these components create a workflow where narrative text, visual artifacts, and organizational structures remain continuously aligned, enabling writers to externalize, manipulate, and recombine ideas and materials across modalities.
@@ 4.1. Key Features
Figure 2. Vistoria’s interface: a left text editor, a central Cluster panel(can be collapsed when not used), and a right free-form Canvas. Cards (image-text pairs with main object keywords) are generated from user intentions and serve as the basic units; two modes—Exact Craft and Creative Spark—support precise realization vs. exploratory variation.
@@ 4.1.1. Reifying Intention through Multimodal Generation
Writers often externalize early intentions through sketches, short notes, or reference images, but these expressions are typically fleeting and difficult to refine into coherent narrative materials. Supporting the preservation and refinement of such multimodal intentions is therefore essential for sustaining ideation (DG1). To address this, Vistoria reifies multimodal inputs into persistent Cards that bundle an image, a story fragment, and extracted narrative objects (characters, objects, scenes). Drawing on the principle of Reuse, Vistoria reifies sketches and text from ephemeral marks into manipulable and reusable artifacts that remain on the canvas and can be invoked again as elements for regeneration, enabling writers to carry forward and iteratively develop their creative intentions. At the same time, we treat cards as alignable units: text and their corresponding images are designed to carry the same underlying meaning, with images serving as visualizations of the narrative described in text.
@@
Vistoria further balances precision and exploration by offering two complementary generation modes (Figure 3). In Exact Craft mode, single cards closely adhere to the author’s expressed intention to concretize specific ideas. In Creative Spark mode, three cards were generate represent diverse options based on writer’s intention, the system deliberately introduces variation around characters, settings, or objects, providing alternative prompts that can inspire new directions.
Figure 3. Vistoria balances precision and exploration through two complementary generation modes: in Exact Craft mode, a single card closely follows the author’s intention to concretize ideas; in Creative Spark mode, three cards are generated with deliberate variations in characters, settings, or objects, offering diverse options to inspire new directions.
Figure 4. A set of visual instruments for image-text co-editing to enhance planning and translating of fictional story writing: (a) Lasso selects regions for coupled image–text edits. (b) Writers extract elements and compose across cards to discover new narrative directions, using collage alongside other instruments to recombine sketches, images, and text on the canvas. (c) Perspective Shift changes an image’s viewpoint and automatically regenerates the story’s point of view (first/third/second person). (d) Filters align visual style and textual tone (e.g., melancholic/dreamy) by jointly altering image effects and rewriting prose.
4.1.2. Synchronized Image–Text Co-Editing through Visual Instruments
Fictional story writing benefits from fluid movement between abstract textual reasoning and concrete visual imagination, yet existing tools separate these modalities and disrupt creative flow (DG2, DG3). To address this, we introduce a set of visual instruments (Figure 4) designed around three principles: (1) Reification, which draws on familiar operations from image editing (e.g., lassoing, collaging, filtering) and transforms them into persistent, manipulable operations that apply equally to text and image elements (Swerzenski, 2021; Masson et al., 2024a); (2) Polymorphism, which ensures the same operations apply uniformly across text and images; and (3) Structural Mapping, which synchronizes changes so that visual and textual representations carry the same narrative meaning. Together, these principles reduce modality switching costs while maintaining coherence between story segments and imagery.
@@ 4.1.2. Synchronized Image–Text Co-Editing through Visual Instruments
Lasso. The lasso instrument exemplifies reification by turning the abstract action of “focusing on part of a story” into a manipulable unit: selecting a region in either an image or a fragment of text generates a new card focusing on the selected part with enriched narrative and visual details. Through polymorphism, the same selection logic applies across modalities—whether circling a visual detail or isolating a text segment—providing a consistent interaction pattern. The text within the selected area in the original content will be emphasized to form a new card. The extracted portion of text is used to regenerate the corresponding image. Structural mapping ensures that the lassoed image region and the expanded story fragment correspond to one another, aligning visual and textual perspectives within the same narrative unit (Figure 4 (a) ).
@@
Collage. The collage instrument reifies the abstract act of “recombining inspirations” into a tangible manipulation: fragments of images, sketches, or text can be directly composed within a collage frame to form a new card. Through polymorphism, the same cut–paste–combine logic applies uniformly across modalities—an image region, a text excerpt, or a sketch element can all be treated as compositional materials for intention-based generation. Guided by structural mapping, the system interprets the spatial arrangement of these multimodal pieces as narrative intent, generating a card where textual descriptions and visual depictions are aligned. For instance, merging two character fragments not only produces a combined image but also generates a new story segment situating them together, ensuring that narrative and imagery evolve in sync (Figure 4 (b)).
@@
Filter. Stylistic coherence is critical in fictional story writing, as consistent affective and aesthetic cues sustain narrative transportation (Green and Brock, 2000), activate readers’ interpretive schemas (Arbib, 1992), and enhance the emotional resonance of literariness (Miall and Kuiken, 1999). In Vistoria, the filter instrument reifies this abstraction into a concrete tool: applying a “melancholic” or “dreamy” filter adjusts both the style of the image and the tone of the accompanying prose (Appendix Table 4). Through polymorphism and structural mapping, the same filter operation works seamlessly across modalities, leveraging the correspondence between visual style in images and emotional tone in text to act on both simultaneously. By making intangible stylistic intentions manipulable and synchronized, filters expand expressive possibilities while maintaining narrative immersion (Figure 4 (c)).
@@
Perspective Shift. Fictional story writing often utilizes perspective shifts, as narratology highlights that changes in voice and focalization fundamentally reshape how events and characters are perceived (Genette, 1980). Cognitive poetics further shows that such shifts alter readers’ empathy and immersion. First-person narrations foster intimacy, while third-person perspectives enable broader structural awareness (Keen, 2007). The perspective-shift instrument reifies this narratological concept into an actionable operation: changing the visual viewpoint of a scene automatically regenerates the story fragment from a first-, third-, or second-person perspective. Through polymorphism, this instrument applies consistently across modalities, altering either an image or its accompanying text triggers a corresponding adjustment in the other. Structural mapping ensures that the focalization shift carries the same meaning across text and image: a new camera angle in the image corresponds to a new narrative voice in the text, allowing writers to explore empathy, distance, and structural awareness in a synchronized manner (Figure 4 (d)).
@@
On the canvas, writers can highlight textual segments, edit stories directly, and add inline comments noting potential uses in later drafting. Story objects, such as characters, settings, or scenes, are represented as editable keywords that can be highlighted by themselves. The system automatically binds each highlighted object to its associated text, consolidating references across multiple cards.
The cluster panel then aggregates all highlighted objects into an organized overview of evolving narrative elements. This eliminates the need to manually scan scattered cards and provides writers with a dynamically updated, object-centered workspace. Selecting an object reveals its complete associated materials, such as linked images, highlighted text segments, and comments, which creates a multimodal, context-rich reference for downstream writing.
Beyond simple aggregation, the panel supports higher-level knowledge construction through its summary feature. When this feature is invoked, the system generates structured summaries of settings, descriptions, and plot elements derived from highlights and comments. These summaries distill fragmented annotations into narrative building blocks (Burroway et al., 2022), enabling writers to iteratively scaffold coherent storylines from previously disjointed ideas.
@@ 4.2. Implementation
System Architecture. We adopted a decoupled front-end/back-end architecture. The React front-end enables efficient rendering for complex interactive interfaces, while the Flask back-end flexibly handles model calls with minimal overhead. Axios manages asynchronous communication between layers. The canvas module uses React-Flow for dynamic node-edge relationships and integrates “path to SVG” for natural interactions like freehand drawing and lasso selection. A DOM-based screenshot function ensures visual consistency in exported images. The text editor integrates React-Quill for rich text editing with lightweight customizations. Global state (text content, canvas nodes, selections) is centrally managed via Zustand to ensure consistency across modules. Per-session persistence relies on sessionStorage, which preserves data during the study session and clears on tab close to protect privacy.
Back-end Multi-Agent flow. The back-end bridges front-end actions with language and vision models through multi-agent orchestration. Every canvas action (e.g., selection, lasso, collage, inline text) is rasterized into screenshots paired with local context (outline, prior cards, global theme). A text agent (o4-mini) examines screenshots using templated prompts from a prompt library to extract user intent, recover sketch layouts, and write text consistent with surrounding outlines. The agent returns structured JSON (story text, intent, layout hints, semantic tags) that downstream visual agents use for synthesis and alignment.
Image Generation and Editing. An image agent invokes Flux Kontext Pro with consolidated prompts that fuse screenshot-derived intent, story fragments, and style controls. Sketches serve as structural scaffolds respecting composition while avoiding duplication; reference images act as content anchors, encouraging variation over copying. When neither is present, Flux operates in free mode guided by text-agent intent and style.For editing, front-end instruments (e.g., perspective shifts, filters, collage) are translated into prompt deltas through GPT-4o’s image pathway, yielding coupled visual updates. In parallel, the text agent revises tone or viewpoint for synchronized multimodal transformation. All artifacts are persisted as cards (image, story, objects keywords), enabling highlight-and-cluster operations that help reassemble fragments into coherent narratives.
@@ 5. User Study
To evaluate the usability, effectivness and usefulness of Vistoria, we conduct a lab user study with 12 participants. The study aims to answer the following research questions:
  • RQ1: What interaction patterns and strategies emerge when participants use a image-text co-editing process?
  • RQ2: How does image-text co-editing facilitate and shape the cognitive process of story creation, in terms of ideation, iteration, and refinement?
@@ 5.1. Participants
We recruited 12 participants (6 males, 6 females, aged 21–32, M=25.5), all with prior creative writing experience. Most held Bachelor’s or Master’s degrees, with backgrounds spanning science, arts, design, or communication. Their creative practices included fiction writing, screenwriting, songwriting, advertising, research, and philosophy. Participants’ creative writing experience ranged from under one year (n=5) to over seven years (n=1), with others reporting 1–3 years (n=2) or 4–7 years (n=3).
All participants were familiar with large language models (e.g., ChatGPT, Gemini, Claude) and had used them for idea generation, editing, descriptive support, content expansion, worldbuilding, and style imitation. This diversity reflects our design goal of supporting not only expert writers but also a broader range of users. Each participant received \$40 USD compensation.

### v3

@@ 4. Vistoria System
In this section, we present the key features of Vistoria. As shown in Figure 3, the interface comprises three primary components: a left Text Editor, a central collapsible Cluster panel, and a right Canvas interface. The Text Editor displays the current story draft, serving as contextual information for content generation. The right Canvas supports freeform sketching, text input, and image-text generation and editing tools. The central Cluster panel aggregates highlights and annotations from Canvas, displaying related plots, settings, and descriptions of each highlighted element for easy reference and overview.
Figure 4. Vistoria analyzes selected text, sketches, images, and broader story context to generate image–text cards that externalize narrative development directions.
The figure shows how Vistoria generates image–text cards from multimodal input. Selected regions containing text, sketches, and images on the Canvas are analyzed together with the broader story context. The system then produces image–text cards that externalize potential narrative development directions.
Figure 5. A set of instrumental operations for image-text co-editing to enhance planning and translating of fictional story writing: (a) Lasso selects regions for coupled image–text edits. (b) Collage enables writers to extract elements and compose across cards to discover new narrative directions. (c) Perspective Shift changes an image’s viewpoint and automatically regenerates the story’s point of view (first/third/second person). (d) Filters align visual style and textual tone (e.g., melancholic/dreamy) by jointly altering image effects and rewriting prose.
The figure illustrates four instrumental operations for synchronized image–text co-editing. The Lasso operation selects localized regions for focused edits, the Collage operation combines extracted elements across cards, the Filter operation aligns visual style with textual tone, and the Perspective-Shift operation changes visual viewpoint and narrative voice together. Each operation affects images and text simultaneously.
@@ 4.1.1. Reifying Intention through Multimodal Generation
The system enables writers to externalize early, vague ideas using multimodal inputs (DG1). To support this, Vistoria transforms multimodal input (sketches, text, and images) into cards that pair a representative image with its corresponding narrative segment. Writers select a region of interest, capturing all multimodal elements within that area. Vistoria then infers from these multimodal elements and broader story context to generate image–text cards that reflect the writer’s intention for narrative direction (Figure 4). These cards serve a dual purpose: they transform ephemeral multimodal input into persistent, regenerable artifacts on the Canvas, while also functioning as semantically aligned units where image and text convey the same underlying meaning.
@@
Vistoria further balances precision and exploration by offering two complementary generation modes. In Exact Craft mode, single cards closely adhere to the author’s expressed intention to concretize specific ideas. In Creative Spark mode, three cards are generated to represent diverse options based on the writer’s intention. The system deliberately introduces variation around characters, settings, or objects, providing alternative prompts that can inspire new directions.
The figure shows how highlighted text segments and objects on cards are aggregated in the Cluster panel. Narrative elements are grouped by character, object, or scene, and the system generates structured summaries of plots, settings, or descriptions. Selecting an item reveals its associated images, text highlights, and comments to support coherent story construction.
4.1.2. Synchronized Image–Text Co-Editing through Instrumental Operations
Fictional story writing benefits from fluid movement between abstract textual reasoning and concrete visual imagination. The system should tightly align text and images so that edits in one modality fluidly inform the other (DG2, DG3). To address this, we introduce a set of Instrumental Operations (Figure 5) designed around three principles: (1) Reification (instrumental interaction), which draws on familiar image-editing operations to make abstract image–text co-editing actions more concrete and manipulable; (2) Polymorphism (instrumental interaction), designing a set of instrumental operations (Lasso, Collage, Filter, Perspective Shift) which ensure the same operations apply uniformly across text and images to lower switching cost; and (3) Dual Coding Theory, which indicates that verbal and nonverbal changes should be aligned to maintain coherence across cognitive channels.
Lasso.
@@ 4.1.2. Synchronized Image–Text Co-Editing through Instrumental Operations
The Lasso instrument exemplifies reification by turning the abstract action of “focusing on part of a story” into a manipulable unit: selecting a region in either an image or a fragment of text triggers the generation of a new card focusing on the selected part with enriched narrative and visual details. Through polymorphism, the same selection logic applies across modalities—whether circling a visual detail or isolating a text segment—providing a consistent interaction pattern. The text within the selected area in the original content will be emphasized to form a new card. The extracted portion of text is used to regenerate the corresponding image. The lassoed image region and the expanded story fragment correspond to one another, aligning visual and textual perspectives within the same narrative unit (Figure 5 (a)).
@@
Collage. The Collage instrument reifies the abstract act of “recombining inspirations” into a tangible manipulation: fragments of images, sketches, or text can be directly composed within a collage frame to form a new card. The same cut–paste–combine logic applies uniformly across modalities—an image region, a text excerpt, or a sketch element can all be treated as compositional materials for intention-based generation. The system interprets the spatial arrangement of these multimodal pieces as narrative intent, generating a card where textual descriptions and visual depictions are aligned. For instance, merging two character fragments not only produces a combined image but also generates a new story segment situating them together, ensuring that narrative and imagery evolve in sync (Figure 5 (b)).
@@
Filter. Stylistic coherence is critical in fictional story writing, as consistent affective and aesthetic cues sustain narrative transportation (Green and Brock, 2000), activate readers’ interpretive schemas (Arbib, 1992), and enhance the emotional resonance of literariness (Miall and Kuiken, 1999). In Vistoria, the Filter instrument reifies this abstraction into a concrete tool: applying a “melancholic” or “dreamy” filter adjusts the visual style and rewrites the accompanying prose to match the emotional tone (Appendix Table 3). Through polymorphism, the same filter operation works seamlessly across modalities, leveraging the correspondence between visual style in images and emotional tone in text to simultaneously act on both. By making intangible stylistic intentions manipulable and synchronized, filters expand expressive possibilities while maintaining narrative immersion (Figure 5 (c)).
@@
Perspective Shift. Fictional story writing often utilizes perspective shift, and narratology highlights that changes in voice and focalization fundamentally reshape how events and characters are perceived (Genette, 1980). Cognitive poetics further shows that such shifts alter readers’ empathy and immersion. First-person narrations foster intimacy, while third-person perspectives enable broader structural awareness (Keen, 2007). The Perspective-Shift instrument reifies this narratological concept into an actionable operation: changing the visual viewpoint of a scene automatically regenerates the story fragment from a first-, third-, or second-person perspective. Through polymorphism, this instrument applies consistently across modalities, altering either an image or its accompanying text triggers a corresponding adjustment in the other. The shift carries the same meaning across text and image: a new camera angle in the image corresponds to a new narrative voice in the text, allowing writers to explore empathy, distance, and awareness in a synchronized manner (Figure 5 (d)).
@@
On the Canvas, writers can highlight textual segments, edit stories directly, and add inline comments noting potential uses in later drafting. Story objects, such as characters, settings, or scenes, are represented as editable keywords that can be highlighted by themselves. The system automatically links each highlighted object to its associated text, consolidating references across multiple cards.
The Cluster panel then aggregates all highlighted objects into an organized overview of evolving narrative elements. This eliminates the need for manual scanning of scattered cards and provides writers with a dynamically updated, object-centered workspace. Selecting an object reveals its complete set of associated materials, including linked images, highlighted text segments, and comments, which creates a multimodal, context-rich reference for downstream writing. Beyond simple aggregation, the panel supports higher-level knowledge construction through its summary feature. When this feature is invoked, the system generates structured summaries of settings, descriptions, and plot elements derived from highlights and comments. These summaries distill fragmented annotations into narrative building blocks (Burroway et al., 2019), enabling writers to iteratively scaffold coherent storylines from previously disjointed ideas.
@@ 4.2. Implementation
We adopted a decoupled front-end/back-end architecture. The React [^2] https://react.dev/ front-end enables efficient rendering for complex interactive interfaces, while the Flask [^3] https://flask.palletsprojects.com/ back-end flexibly handles model calls. Axios manages asynchronous communication between layers.
4.2.1. Front-end.
The front-end consists of three main modules: the Canvas, the Cluster, and the Text Editor. Zustand [^4] https://github.com/pmndrs/zustand centrally manages the global state (including canvas nodes, cluster selections, and text content) to ensure consistency across all modules. To protect privacy during user studies, per-session data is stored in sessionStorage and automatically clears when the tab closes, while users can manually export Canvas nodes and text content via the top toolbar.
Canvas Module. The right-side Canvas module consists of four distinct layers: (1). Node Interaction Layer (Bottom): This layer uses React-Flow [^5] https://reactflow.dev/ to maintain a dynamic node-edge graph. Node types include card, collage, text, sketch, handwriting, and image. All nodes share basic properties, such as ID, type, and coordinates, for edge linking, but the internal data structures of those node types vary for rendering. For example, the “card” node includes features such as image modification tools, image lasso selection, image highlighting, object manipulation, and basic information display. (2). Pen-based Input Capture Layer: This layer supports natural interactions like freehand sketching, writing, and lasso selection. It uses the perfect-freehand library [^6] https://github.com/steveruizok/perfect-freehand to smooth captured points and convert them into Scalable Vector Graphics (SVGs). Each captured SVG stroke is added to the graph as a new node. (3). Generation Selection Layer: A Document Object Model (DOM) based screenshot function ensures visual and positional consistency. After clicking generate, the front-end sends both the screenshot and structured data of the nodes inside the selected area to the backend, adds a new node to the graph, and awaits the returned data. (4). Tool Layer (Top): This is the most visible layer. It contains operation tools and Canvas control tools.
Cluster Module. The middle module displays selected information, including objects, images, text, and annotations. Users filter this content from the node-edge graph by selecting specific objects. A button allows users to expand or collapse this entire area.
Text Editor Module. The left-side Text Editor uses the React-Quill [^7] https://github.com/zenoamaro/react-quill component. It provides lightweight rich-text editing capabilities designed to align with the narrative structure.
4.2.2. Back-End Multi-Agent Flow.
The backend is organized as a multi-agent pipeline, where each prompt-specialized LLM agent is organized sequentially. Rather than relying on a single model, the system decomposes the workflow into three cooperating functional agents:
Narrative Construction Agent. This agent takes multimodal context—such as Canvas screenshots which incorporate all information (sketches, text inputs, and images) contained within the user-selected region to produce structured textual outputs—including user intentions and story segments. Its prompts enforce consistency with contextual information such as the existing story content in the Text Editor and the global stylistic constraints. In essence, the agent transforms the user’s multimodal inputs into coherent narrative storylines. During implementation, the GPT-o4-mini [^8] We used OpenAI’s o4-mini model via the OpenAI API (model ID o4-mini) in August 2025 (OpenAI, 2025). is used to process Canvas screenshots and contextual information to infer user intentions because of its rapid inference speed and strong reasoning ability. GPT-4o [^9] We used OpenAI’s GPT-4o model via the OpenAI API (model ID gpt-4o) in August 2025 (OpenAI, 2024). is used to refine these inferred intentions from o4-mini and contextual information into polished, coherent story segments. Prompts for precise description generation of o4-mini are shown in Appendix A.5). When applying instrumental operations for editing, the story segment from the previous card is also read in prompts and modified accordingly.
Visual Synthesis Agent. Using the narrative produced by the Narrative Construction Agent as prompts, this agent supports image generation using either the GPT-4o API or the FLUX diffusion model [^10] We used the FLUX.1 diffusion model via the Black Forest Labs API in August 2025 (Black Forest Labs, 2024).. When reference images or screenshots are available—such as during instrumental-operation edits that modify images or when multimodal inputs include a base image or sketches—GPT-4o is used for image generation, leveraging its strong capabilities in image understanding and re-generating based on base images. In all other cases, the FLUX diffusion model is used due to its faster generation speed.
Memory Agent. The backend maintains a persistent, globally accessible memory of all previously generated image-text pairs. This agent coordinates read/write operations to this store, enabling multi-turn reuse, cross-scene integration, and contextual continuity. When the user later requests to merge scenes, change details, or perform local edits, this agent retrieves relevant image-text pairs and passes them back to the earlier agents to re-initiate the pipeline.
@@ 5. User Study
To evaluate the usability of Vistoria and to understand how these multimodal interactions design support creativity, we conducted an exploratory lab user study with 12 participants. We structure our user study around two complementary components:
(1) A usability evaluation of Vistoria aiming to answer the following research questions:
RQ1: How useful are the multimodal co-editing functions, and in what ways do they influence participants’ workflows?
RQ2: How does multimodal image–text co-editing affect participants’ workload compared to a text-only baseline?
(2) A pilot study examining the creativity support provided by Vistoria, focusing on:
RQ3: How does multimodal image–text co-editing influence participants’ ideation and the development of fictional stories?
RQ4: How does multimodal image–text co-editing influence participants’ sense of agency and ownership?
RQ1 evaluates the usefulness of these multimodality functions and the strategies participants adopted in relation to DG3 (polymorphic instrumental operations). RQ3 explores how DG1 (reifying intentions through multimodal input) and DG2 (image-text alignment) support creative ideation and narrative development. The design of the cluster panel corresponding to DG4 functions mainly as auxiliary support and is not central to our user evaluation of multimodal interaction.
Note that we used the sense of ownership to refer to the writer’s “sense of possession” over the resulting narratives within the system, even the AI-generated artifacts (He et al., 2025; Lee et al., 2024; Draxler et al., 2024b; Weber et al., 2024). The declaration of the sense of agency, on the other hand, refers to the writer’s awareness of “initiating, executing, and controlling” key actions in the writing and artifact editing process (Moruzzi, 2022; Legaspi et al., 2019; Lee et al., 2024).
@@ 5.1. Participants
As the system targets intermediate to expert participants who understand narrative structure and already use LLMs in their creative workflows, we recruited 12 participants (6 males, 6 females, aged 21–32, M=25.5), all with prior creative writing experience. Participants were also recruited through student organizations by sharing our study announcement in their group chats. All held Bachelor’s degrees, with backgrounds spanning science, arts, design, or communication. Their creative practices included fiction writing, screenwriting, songwriting, advertising, and media production. Participants’ creative writing experience ranged from under one year (n=5) to over seven years (n=1), with others reporting 1–3 years (n=3) or 4–7 years (n=3). All participants were familiar with LLMs (e.g., ChatGPT, Gemini, Claude) and had used them for idea generation, editing, descriptive support, content expansion, world-building, and style imitation. Among the 12 participants, 6 use LLMs daily, while the remaining 6 are evenly split across several times a week, occasionally, and rarely (2 in each). Each participant received a \$40 USD compensation after finishing the experiment.